# A data-driven analysis of the impact of non-compliant individuals on epidemic diffusion in urban settings

### v2

@@ 1 Introduction
Over the past decade, considerable research has focused on understanding how individual behavior influences the dynamics of epidemic outbreaks, demonstrating that behavioral factors can substantially alter the course and severity of disease spread [1, 2, 3, 4, 5]. During the COVID-19 pandemic, unsafe health behaviors were often associated with the spread of misinformation on online social networks [6, 7, 8]. Health-related non-compliance encompasses a range of risk-enhancing behaviors, including increased vaccine refusal, adoption of unproven or harmful ”treatments”, and the disregard of public health measures such as social distancing and mask mandates [8, 9, 10]. Vaccine hesitancy in particular has been linked to the resurgence of vaccine-preventable diseases such as measles, previously believed to be eliminated in many regions [11, 12]. Collectively, such behaviors pose significant challenges during epidemic events, as they may amplify transmission rates and exacerbate the overall impact of outbreaks [13].
To investigate the impact of individuals not complying with public health safety measures on epidemic dynamics, several studies have extended agent-based Susceptible-–Infected-–Recovered (SIR) models [14] by allowing disease transmission to depend on individual behavioral status [4, 2, 15, 3]. In these models, individuals may isolate themselves in response to external mandates (e.g., lockdowns) or perceived risk, thereby effectively removing themselves from the transmission network. More recently, DeVerna et al. [13] modeled the impact of non-compliant individuals through large-scale data-driven simulations of contact networks combined with measurements of online misinformation consumption on social media. Their model increases the infectiousness (i.e., the probability of transmitting the disease) of ”misinformed” individuals, but keeps susceptibility constant across the population, regardless of behavioral status. However, earlier research has shown that heterogeneity in both infectivity and susceptibility plays a critical role in shaping epidemic outcomes on contact networks [16, 17, 18]. To fully capture the effects of behavioral differences, models should account for variation in both dimensions.
@@
To explore the impact of non-compliant individuals on epidemic dynamics in urban environments, we construct detailed contact networks that capture in-person interactions among city residents. These networks are generated using a data-driven approach that integrates the geographic distribution of the population with age-specific contact matrices derived from survey data [22]. The resulting age-stratified contact graphs assign individuals to localized spatial units (tiles) and reflect realistic patterns of interaction. We focus on three major Italian cities—Torino, Milano, and Palermo—with populations ranging from approximately 600,000 to 1.2 million residents. To capture behavioral heterogeneity, we employ the HeSIR model [17], a heterogeneous extension of the classic SIR framework [14], which differentiates between ordinary individuals (O) and non-compliant individuals (M, short for Misbehaving). The latter group represents those who disregard public health recommendations, such as mask-wearing, social distancing, or vaccination. Their behavior increases both their susceptibility to infection and their ability to infect others, making them a key risk factor in disease propagation. Using efficient epidemic simulation techniques, we quantify how the presence of non-compliant individuals affects key epidemic indicators, examining sensitivity to both model parameters and the proportion of such individuals. To incorporate spatial heterogeneity in behavior, we introduce a tile-level non-compliance propensity score, derived from electoral data and studies linking political orientation to vaccine hesitancy [23, 24]. This produces a data-driven spatial distribution of non-compliance within each city. Our results show that even small proportions of non-compliant individuals can substantially alter epidemic outcomes. Moreover, when these individuals are geographically clustered, they can create local infection hotspots that may overwhelm nearby healthcare services. These findings underscore the critical importance of monitoring behavioral compliance within urban areas and tailoring public health interventions to address localized risks.
@@ 2.1 Contact graph
City Tile size $N$ $N_{T}$ $N_{H}$ $k\right>\pm\sigma_{k}$ $N_{i}\right>\pm\sigma_{N_{i}}$
@@
Table 1: Summary statistics of the contact graphs generated for the cities under analysis. $N$ denotes the number of individuals; $N_{T}$ is the number of spatial tiles; and $N_{H}$ is the number of households. The average degree of the graph is given by $k\right>$, and the average number of individuals per tile by $N_{i}\right>$. The standard deviations of the degree and the number of individuals per tile are denoted by $\sigma_{k}$ and $\sigma_{N_{i}}$, respectively.
@@ 2.2 HeSIR epidemic model
To study the effect of individuals who do not comply with public health recommendations, hereby non-compliant individuals, we employ the HeSIR (Heterogeneous SIR[17] epidemic model, a modified version of the traditional Susceptible–Infectious–Recovered (SIR) model [14].
@@ 2.3 Data-driven distribution of non-compliant individuals
In our work, we make use of these coefficients under the hypothesis that vaccine hesitancy serves as a proxy for individual non-compliance with public health directives. For each geographic tile $i$, we define the proportion of non-compliant individuals, denoted $r_{i}$, which is rescaled to the interval $[0,1]$—where 00 corresponds to minimal non-compliance and $1$ to maximal non-compliance—independently for each city. To compute the values of $r_{i}$ and thereby obtain the spatial distribution of the non-compliant population, we combine the VHE fit coefficients with electoral results.
The electoral data, originally reported at the level of electoral districts, are projected onto the tiles proportionally to their area of overlap. Mathematically, this is expressed as
@@
Figure 3: Data-driven distribution of the proportion of non-compliant individuals, $r_{i}$, scaled such that 00 represents the minimum level and $1$ the maximum level of non-compliance. Darker colors correspond to smaller values, using the same color normalization for the three cities.
@@ 3.1 The impact of non-compliant individuals: SIR vs HeSIR with uniform distribution
Figure 4 (top panels) shows that when $a,b>1$—representing a scenario where non-compliant individuals are both more infectious and more susceptible than the rest of the population—the total number of infected individuals increases significantly compared to the baseline SIR model, for all values of $\beta$. These effects are quantified by the Attack Rate (AR), $R_{\infty}$, defined as the total number of individuals recovered at the end of the outbreak. We also observe that below a certain threshold value of $\beta$, the epidemic fails to spread, resulting in $R_{\infty}\approx 0$. This behavior is a well-known characteristic of the SIR model. Notably, the critical transmission rate $\beta_{c}$ required for an outbreak decreases as $a$ and $b$ increase, and to a lesser extent as the fraction of non-compliant individuals $p_{M}$ increases. In other words, the presence of non-compliant individuals lowers the transmission rate threshold $\beta$ needed for the disease to infect a significant portion of the population, especially when non-compliant individuals exhibit substantially higher infectivity and susceptibility (larger $a$ and $b$).
The bottom panels of Figure 4 display the difference in Attack Rate (AR) between the HeSIR model with a uniform distribution of non-compliant individuals and the baseline SIR model. The increase in AR is striking, peaking around $\beta\langle k\rangle=0.07$, where an additional 40% of the population becomes infected for $p_{M}=0.3$. This peak aligns with the epidemic threshold of the SIR model, at which $R_{\infty}\approx 0$. However, the AR difference remains substantial even at intermediate values of $\beta$. For example, at $\beta\langle k\rangle=0.1$, the additional fraction of infected individuals ranges from 5% to 15% for $p_{M}=0.1$, from approximately 10% to 25% for $p_{M}=0.2$, and from about 12% to 33% for $p_{M}=0.3$, with the highest values corresponding to the riskiest behaviors (larger $a,b$). Remarkably, the extra AR decreases rapidly at very high transmission rates, indicating that the presence of non-compliant individuals has a more pronounced impact when disease transmissibility is moderate.
@@
Figure 4: Simulations of the HeSIR model in the city of Milano with a uniform distribution of non-compliant individuals. Panels (a-c): total fraction of infected individuals (Attack Rate) as a function of the product of transmission rate $\beta$ and average graph degree $k\right>$, for varying values of $a$, $b$, and $p_{M}$. Panels (d-f): difference in Attack Rates between the HeSIR and baseline SIR models, with the same $p_{M}$ as the panel above. In all panels, lines indicate the mean and shaded areas correspond to inter-quartile range.
@@
Figure 5: Epidemic outbreak simulations of the HeSIR model in the city of Milano with a uniform distribution of non-compliant individuals ($\beta\langle k\rangle=0.1$). Panels (a-b): time evolution of the infected population fraction for two values of the non-compliant fraction $p_{M}$. Lines indicate median values and shaded areas the interquartile range. Panel (c): scatter plot of epidemic peak time versus peak amplitude, averaged over 100 simulations, for various combinations of model parameters. Points located in the upper-left region indicate more severe outbreaks (earlier and higher peaks). The baseline SIR scenario, with no behavioral heterogeneity ($a=b=1$), is shown in grey.
@@ 3.2 The impact of non-compliant individuals: HeSIR with uniform vs data-driven distribution
To quantify the relationship between the local fraction of non-compliant individuals $p_{M,i}$ and the resulting change in attack rate, we perform a linear regression of $\Delta\left<R_{\infty}\right>_{i}$ against $p_{M,i}$, weighting each tile by its population. The fitted lines for selected scenarios are shown in the bottom panels of Fig. 7. We observe that the slope of the fit, which captures the sensitivity of the local outbreak severity to variations in non-compliance, varies markedly with the transmission rate $\beta$. To explore this dependence systematically, we compute the slope of the linear fit across a wide range of parameter combinations $(\beta,a,b)$, with results summarized in Fig. 8 (left panel). Remarkably, all cities exhibit consistent qualitative trends (see Supplemental Material): the slope initially increases with $\beta$, reaches a peak, and subsequently declines toward an asymptotic value. This decline reflects the diminishing marginal effect of behavioral heterogeneity as disease transmissibility increases. When $\beta$ is high, even modest levels of non-compliance suffice to drive widespread infection, reducing the impact of spatial clustering of non-compliant individuals.
@@ 3.3 Effects of contact graph randomization
As detailed in Sec. 2.1, the contact graph model employed in this study captures the complexity of both household and social interactions, with social contacts influenced by factors such as age class, geographic distance, and individual fitness. To evaluate the impact of these modeling assumptions, we construct a randomized version of the contact graph using the configuration model [32], which preserves the original degree sequence but randomly reassigns edges. This shuffled contact graph maintains the total number of connections but effectively eliminates almost all household ties as well as correlations related to age, distance, and individual fitness in social contacts.
@@ 4 Discussion
In this study, we investigated the impact of non-compliant individuals on epidemic propagation within urban environments. Leveraging detailed geographic data, we constructed large-scale contact networks—comprising up to 1.2 million nodes—for three Italian cities, explicitly distinguishing between household-based and broader social interactions. Our minimally modified SIR framework incorporates a subpopulation of non-compliant individuals, which leads to a substantial increase in the total number of infections and a marked acceleration of epidemic dynamics. The resulting amplification of the infection force enables outbreaks to occur even for diseases with lower intrinsic transmission rates. Moreover, we find that the peak number of infections rises significantly when non-compliant behavior is more pronounced, even if such individuals represent only a small fraction of the urban population. These results underscore the critical role that behavioral heterogeneity – specifically, non-compliance – can play in facilitating the spread of infectious diseases.
@@
Building on empirical links between health-related non-compliance and political affiliation, we constructed a data-driven, geographically heterogeneous distribution of non-compliant individuals and compared its epidemiological effects to those of an homogeneous distribution. While city-level epidemic metrics, such as attack rate, peak time, and peak amplitude, showed limited but non-negligible differences between the two scenarios, substantial disparities emerged at finer spatial scales, i.e., at the tile level.
@@
We observed that both the local attack rate and epidemic peak were strongly influenced by the local concentration of non-compliant individuals. As expected, when the spatial distribution of non-compliant individuals is heterogeneous, the epidemic tends to intensify in regions with a higher density of non-compliant individuals and remains comparatively subdued in areas with lower density. Notably, this spatial dependence becomes more pronounced when the disease’s baseline transmission rate is just above the epidemic threshold, indicating that behavioral heterogeneity exerts a greater influence on marginal outbreak conditions.
@@
When comparing these results to those obtained using a shuffled version of the contact graph, we observed only marginal differences at the city-wide scale, but substantial effects at the local (tile) level. In particular, the observed dependence of local attack rates and epidemic peaks on the fraction of non-compliant individuals was significantly reduced for certain values of the misbehavior parameters, and in some cases nearly eliminated. These findings highlight the critical role played by the two-level city contact graphs employed in our study, which maintain a realistic balance between local (household) and long-range (social) connections. Randomly shuffling edges disrupts this structure, altering the ratio of internal to total connections within tiles and thereby modifying the network’s spatial organization. This disruption increases the number of potential transmission pathways, reducing the influence of localized behavioral clustering and diluting the spatial heterogeneity that would otherwise shape epidemic outcomes.
@@
In summary, our results demonstrate that the presence of non-compliant individuals can significantly alter epidemic dynamics, even when they constitute only a small fraction of the overall population. Their spatial concentration within specific urban areas is particularly concerning, as it can generate infection “hotspots” that place disproportionate pressure on local healthcare infrastructure. These findings highlight the potential value of geographically targeted public health interventions, both for epidemic preparedness and for behavior-focused campaigns aimed at increasing adherence to public health guidelines, including but not limited to mask usage, physical distancing, and vaccination.
@@
Our method for estimating the spatial distribution of non-compliance was based on political affiliation data and vaccine hesitancy adaptation coefficients. While this provides a plausible proxy for health-related behavioral resistance, it captures only a subset of the broader phenomenon. Future refinements could incorporate socio-economic and demographic variables to improve the realism of misbehavior propensity models. Additionally, our simulations assumed random initial infections, with no preferential spatial placement of index cases. Alternative scenarios, in which outbreaks originate in areas with high numbers of non-compliant individuals or at geographically critical nodes such as transportation hubs, could yield different dynamics and will be explored in future work. Finally, our model considers behavior as static, whereas real-world individuals often adapt their actions during an epidemic, for example by reducing contacts in response to perceived risk [4]. Incorporating such dynamic behavioral feedback mechanisms remains an important direction for future research.
Acknowledgements

### v3

@@ 1 Introduction
Over the past decades, considerable research has focused on understanding how individual behaviour influences the dynamics of epidemic outbreaks, demonstrating that behavioural factors can substantially alter the course and severity of disease spread [1, 2, 3, 4, 5, 6, 7]. During the COVID-19 pandemic, unsafe health behaviours have been associated with the spread of misinformation on online social networks [8, 9, 10]. Health-related non-compliance encompasses a range of risk-enhancing behaviours, including increased vaccine refusal, adoption of unproven or harmful ”treatments”, and the disregard of public health measures such as social distancing and mask mandates [10, 11, 12]. Vaccine acceptance, in particular, is a topic of great importance for public health [13, 14]. The recent rise in vaccine hesitancy has been linked to the resurgence of vaccine-preventable diseases such as measles, previously believed to be eliminated in many regions [15, 16]. Collectively, non-compliant behaviours pose significant challenges during epidemic events, as they may amplify transmission rates and exacerbate the overall impact of outbreaks [17, 18, 19].
To investigate the impact of individuals not complying with public health safety measures on epidemic dynamics, several studies have extended agent-based Susceptible–Infected–Recovered (SIR) models [20] by allowing disease transmission to depend on individual behavioural status [5, 3, 21, 4, 22]. In these models, individuals may isolate themselves in response to external mandates (e.g., lockdowns) or perceived risk, thereby effectively removing themselves from the transmission network. More recently, Deverna et al. [17] modelled the impact of non-compliant individuals through large-scale data-driven simulations of contact networks combined with measurements of online misinformation consumption on social media. Their model increases the infectiousness (i.e., the probability of transmitting the disease) of ”misinformed” individuals, but keeps susceptibility constant across the population, regardless of behavioural status. However, earlier research has shown that heterogeneity in both infectivity and susceptibility plays a critical role in shaping epidemic outcomes on contact networks [23, 24, 25]. To fully capture the effects of behavioural differences, models should account for variation in both dimensions.
@@
To explore the impact of non-compliant individuals on epidemic dynamics in urban environments, we construct detailed contact networks that capture in-person interactions among city residents. These networks are generated using a data-driven approach that integrates the geographic distribution of the population with age-specific contact matrices derived from survey data [29]. The resulting age-stratified contact graphs assign individuals to localized spatial units (tiles) and reflect realistic patterns of interaction. We focus on three major Italian cities—Torino, Milano, and Palermo—with populations ranging from approximately 600,000 to 1.2 million residents. To capture behavioural heterogeneity, we employ the HeSIR model [25], a heterogeneous extension of the classic SIR framework [20], which differentiates between ordinary individuals (O) and non-compliant individuals (M, short for Misbehaving). The latter group represents those who disregard public health recommendations, such as mask-wearing, social distancing, or vaccination. Their behaviour increases both their susceptibility to infection and their ability to infect others, making them a key risk factor in disease propagation. Using efficient epidemic simulation techniques, we quantify how the presence of non-compliant individuals affects key epidemic indicators, examining sensitivity to both model parameters and the proportion of such individuals. To incorporate spatial heterogeneity in behaviour, we introduce a tile-level non-compliance propensity score, derived from electoral data and studies linking political orientation to vaccine hesitancy [30, 31]. This produces a data-driven spatial distribution of non-compliance within each city. Our results show that even small proportions of non-compliant individuals can substantially alter epidemic outcomes. Moreover, when these individuals are geographically clustered, they can create local infection hotspots that may overwhelm nearby healthcare services. These findings underscore the critical importance of monitoring behavioural compliance within urban areas and tailoring public health interventions to address localized risks.
@@ 2.1 Contact graph
City Tile size NN $N_{T}$ $N_{H}$ $\left<k\right>\pm\sigma_{k}$ $\left<N_{i}\right>\pm\sigma_{N_{i}}$
@@
Table 1: Summary statistics of the contact graphs generated for the cities under analysis. NN denotes the number of individuals; $N_{T}$ is the number of spatial tiles; and $N_{H}$ is the number of households. The average degree of the graph is given by $\left<k\right$>, and the average number of individuals per tile by $\left<N_{i}\right$>. The standard deviations of the degree and the number of individuals per tile are denoted by $\sigma_{k}$ and $\sigma_{N_{i}}$, respectively.
@@ 2.2 HeSIR epidemic model
To study the effect of individuals who do not comply with public health recommendations, hereby non-compliant individuals, we employ the HeSIR (Heterogeneous SIR) epidemic model, a modified version of the traditional Susceptible-Infectious-Recovered (SIR) model [20]. This model was proposed and studied in Ref. [25], where the epidemic threshold was derived analytically for a population with homogeneous mixing and for the case of a network with homogeneous contacts. In this section, we introduce the model in detail and discuss its differences from the classical SIR.
@@
It is worth noting that an alternative, though mathematically equivalent, perspective would be to consider non-compliant behaviour as the default risk level and model compliance as a reduction in infection probability compared to that baseline (e.g. [3, 36]). The rationale of our choice lies in the situation experienced in Italy during the COVID-19 pandemic, with a dominant fraction of compliant individuals (e.g., vaccine coverage was approximately 85% with two doses toward the end of 2021 [37]) which naturally leads to considering non-compliant individuals as exceptions.
@@ 2.3 Data-driven distribution of non-compliant individuals
In our work, we make use of these coefficients under the hypothesis that vaccine hesitancy serves as a proxy for individual non-compliance with public health directives. For each geographic tile $i$ specified in Section 2.1, we define the proportion of non-compliant individuals, denoted $r_{i}$, which is rescaled to the interval [0,1][0,1]—where 0 corresponds to minimal non-compliance and $1$ to maximal non-compliance—independently for each city. To compute the values of $r_{i}$ and thereby obtain the spatial distribution of the non-compliant population, we combine the VHE fit coefficients with electoral results.
Election data are originally reported at the level of electoral districts, which vary in size and shape. To obtain a measure of non-compliance for the tiles, votes are projected proportionally to their area of overlap. Mathematically, this is expressed as
@@
Figure 3: Data-driven distribution of the proportion of non-compliant individuals, $r_{i}$, scaled such that 0 represents the minimum level and $1$ the maximum level of non-compliance. Darker colors correspond to smaller values, using the same color normalization for the three cities.
@@ 3.1 The impact of non-compliant individuals: SIR vs HeSIR with uniform distribution
Figure 4 (top panels) shows that when a,b>1a,b>1—representing a scenario where non-compliant individuals are both more infectious and more susceptible than the rest of the population—the total number of infected individuals increases significantly compared to the baseline SIR model, for all values of $\beta$. These effects are quantified by the Attack Rate (AR), $R_{\infty}$, defined as the total number of individuals recovered at the end of the outbreak. We also observe that below a certain threshold value of $\beta$, the epidemic fails to spread, resulting in $R_{\infty}\approx 0$. This behaviour is a well-known characteristic of the SIR model. Notably, the critical transmission rate $\beta_{c}$ required for an outbreak decreases as $a$ and $b$ increase, and to a lesser extent as the fraction of non-compliant individuals $p_{M}$ increases. In other words, the presence of non-compliant individuals lowers the transmission rate threshold $\beta$ needed for the disease to infect a significant portion of the population, especially when non-compliant individuals exhibit substantially higher infectivity and susceptibility (larger $a$ and $b$).
The bottom panels of Figure 4 display the difference in Attack Rate (AR) between the HeSIR model with a uniform distribution of non-compliant individuals and the baseline SIR model. The increase in AR is striking, peaking around $\beta\langle k\rangle=0.07$, where an additional 40% of the population becomes infected for $p_{M}=0.3$. This peak aligns with the epidemic threshold of the SIR model, at which $R_{\infty}\approx 0$. However, the AR difference remains substantial even at intermediate values of $\beta$. For example, at $\beta\langle k\rangle=0.1$, the additional fraction of infected individuals ranges from 5% to 15% for $p_{M}=0.1$, from approximately 10% to 25% for $p_{M}=0.2$, and from about 12% to 33% for $p_{M}=0.3$, with the highest values corresponding to the riskiest behaviours (larger $a,b$). Remarkably, the extra AR decreases rapidly at very high transmission rates, indicating that the presence of non-compliant individuals has a more pronounced impact when disease transmissibility is moderate.
@@
Figure 4: Simulations of the HeSIR model in the city of Milano with a uniform distribution of non-compliant individuals. Panels (a-c): total fraction of infected individuals (Attack Rate) as a function of the product of transmission rate $\beta$ and average graph degree $\left<k\right$>, for varying values of $a$, $b$, and $p_{M}$. Panels (d-f): difference in Attack Rates between the HeSIR and baseline SIR models, with the same $p_{M}$ as the panel above. In all panels, lines indicate the mean and shaded areas correspond to inter-quartile range.
@@
Figure 5: Epidemic outbreak simulations of the HeSIR model in the city of Milano with a uniform distribution of non-compliant individuals ($\beta\langle k\rangle=0.1$). Panels (a-b): time evolution of the infected population fraction for two values of the non-compliant fraction $p_{M}$. Lines indicate median values and shaded areas the interquartile range. Panel (c): scatter plot of epidemic peak time versus peak amplitude, averaged over 100 simulations, for various combinations of model parameters. Points located in the upper-left region indicate more severe outbreaks (earlier and higher peaks). The baseline SIR scenario, with no behavioural heterogeneity ($a=b=1$), is shown in grey.
@@ 3.2 The impact of non-compliant individuals: HeSIR with uniform vs data-driven distribution
To quantify the relationship between the local fraction of non-compliant individuals $p_{M,i}$ and the resulting change in attack rate, we perform a linear regression of $\Delta\left<R_{\infty}\right>_{i}$ against $p_{M,i}$, weighting each tile by its population. The fitted lines for selected scenarios are shown in the bottom panels of Fig. 7. We observe that the slope of the fit, which captures the sensitivity of the local outbreak severity to variations in non-compliance, varies markedly with the transmission rate $\beta$. To explore this dependence systematically, we compute the slope of the linear fit across a wide range of parameter combinations $(\beta,a,b)$, with results summarized in Fig. 8 (left panel). Remarkably, all cities exhibit consistent qualitative trends (see Supplemental Material): the slope initially increases with $\beta$, reaches a peak, and subsequently declines toward an asymptotic value. This decline reflects the diminishing marginal effect of behavioural heterogeneity as disease transmissibility increases. When $\beta$ is high, even modest levels of non-compliance suffice to drive widespread infection, reducing the impact of spatial clustering of non-compliant individuals.
@@ 3.3 Effects of contact graph randomization
As detailed in Sec. 2.1, the contact graph model employed in this study captures the complexity of both household and social interactions, with social contacts influenced by factors such as age class, geographic distance, and individual fitness. To evaluate the impact of these modelling assumptions, we construct a randomized version of the contact graph using the configuration model [41], which preserves the original degree sequence but randomly reassigns edges. This shuffled contact graph maintains the total number of connections but effectively eliminates almost all household ties as well as correlations related to age, distance, and individual fitness in social contacts.
@@ 4 Discussion
In this study, we investigated the impact of non-compliant individuals on epidemic propagation within urban environments. Leveraging detailed geographic data, we constructed large-scale contact networks—comprising up to 1.2 million nodes—for three Italian cities, explicitly distinguishing between household-based and broader social interactions. Our minimally modified SIR framework incorporates a subpopulation of non-compliant individuals, which leads to a substantial increase in the total number of infections and a marked acceleration of epidemic dynamics. The resulting amplification of the infection force enables outbreaks to occur even for diseases with lower intrinsic transmission rates. Moreover, we find that the peak number of infections rises significantly when non-compliant behaviour is more pronounced, even if such individuals represent only a small fraction of the urban population. These results underscore the critical role that behavioural heterogeneity—specifically, non-compliance—can play in facilitating the spread of infectious diseases.
@@
Building on empirical links between health-related non-compliance and political affiliation, we constructed a data-driven, geographically heterogeneous distribution of non-compliant individuals and compared its epidemiological effects to those of a homogeneous distribution. While city-level epidemic metrics, such as attack rate, peak time, and peak amplitude, showed limited but non-negligible differences between the two scenarios, substantial disparities emerged at finer spatial scales, i.e., at the tile level.
@@
We observed that both the local attack rate and epidemic peak were strongly influenced by the local concentration of non-compliant individuals. As expected, when the spatial distribution of non-compliant individuals is heterogeneous, the epidemic tends to intensify in regions with a higher density of non-compliant individuals and remains comparatively subdued in areas with lower density. Notably, this spatial dependence becomes more pronounced when the disease’s baseline transmission rate is just above the epidemic threshold, indicating that behavioural heterogeneity exerts a greater influence on marginal outbreak conditions.
@@
When comparing these results to those obtained using a shuffled version of the contact graph, we observed only marginal differences at the city-wide scale, but substantial effects at the local (tile) level. In particular, the observed dependence of local attack rates and epidemic peaks on the fraction of non-compliant individuals was significantly reduced for certain values of the misbehaviour parameters, and in some cases nearly eliminated. These findings highlight the critical role played by the two-level city contact graphs employed in our study, which maintain a realistic balance between local (household) and long-range (social) connections. Randomly shuffling edges disrupts this structure, altering the ratio of internal to total connections within tiles and thereby modifying the network’s spatial organization. This disruption increases the number of potential transmission pathways, reducing the influence of localized behavioural clustering and diluting the spatial heterogeneity that would otherwise shape epidemic outcomes.
@@
In summary, our results demonstrate that the presence of non-compliant individuals can significantly alter epidemic dynamics, even when they constitute only a small fraction of the overall population. Their spatial concentration within specific urban areas is particularly concerning, as it can generate infection “hotspots” that place disproportionate pressure on local healthcare infrastructure. These findings highlight the potential value of geographically targeted public health interventions, both for epidemic preparedness and for behaviour-focused campaigns aimed at increasing adherence to public health guidelines, including but not limited to mask usage, physical distancing, and vaccination.
@@
While our work relies on generating of large contact graphs and describing the state of each agent/individual, a well-known alternative approach to studying urban epidemics is based on metapopulation models, in which individuals are distributed across different areas (“patches“) within the city and can move from one patch to another during the simulation [27]. This approach may involve a lighter computational load, but it relies on the assumption that the population of a patch is well-mixed, meaning that any individual can be infected by any other. More importantly, metapopulation models are better suited to including mobility in the model: however, we did not have access to such data for this study. Ultimately, we chose the modelling approach put forward by [29] as it provides a proven method for building urban social networks from publicly available datasets.
Our method for estimating the spatial distribution of non-compliance was based on political affiliation data and vaccine hesitancy adaptation coefficients. While this provides a plausible proxy for health-related behavioural resistance, it captures only a subset of the broader phenomenon. Future refinements could incorporate socio-economic and demographic variables to improve the realism of misbehaviour propensity models. Additionally, our simulations assumed random initial infections, with no preferential spatial placement of index cases. Alternative scenarios, in which outbreaks originate in areas with high numbers of non-compliant individuals or at geographically critical nodes such as transportation hubs, could yield different dynamics and will be explored in future work. Finally, our model considers behaviour as static, whereas real-world individuals often adapt their actions during an epidemic, for example by reducing contacts in response to perceived risk [5]. Incorporating such dynamic behavioural feedback mechanisms remains an important direction for future research.
Data Access
The Julia code used for the simulations is available at [42], while the non-compliance distributions obtained for the cities are available in the Github repo [43].
Acknowledgments